Mr Jones is now 28 years older than Robbie. Nine years ago he was five times as old as Robbie was then. How old is each?
Answer by FrankM(1040)

J=R+28
(J-9)=5(R-9)
J-9=5R-45
J=5R-36
R+28=5R-36
64=4R
R=16 J=44